OpenStudy (anonymous):

what is the solution set to -8 through 4

OpenStudy (anonymous):

\[-8\leq x\leq4\]?

OpenStudy (anonymous):

do u even know what a solution set is

OpenStudy (anonymous):

it would be {-7,-6,-5,-4,-3,-2,1,1,2,3,}

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

\[\{x|-8\leq x \leq 4\}\] \[[-8,4]\]

OpenStudy (anonymous):

ssssssssssssssssssssolutions set!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

there should be more values than that

OpenStudy (anonymous):

In mathematics, a solution set is the set of values that satisfy a given set of equations or inequalities. IT DOES NOT HAVE TO BE JUST INTEGER

OpenStudy (anonymous):

so i was right.

OpenStudy (anonymous):

yeah sure,